OpenStudy (anonymous):

Can someone please show me how to solve this? What are the x– and y–intercepts of the graph of y = x2 – 10x + 21?

OpenStudy (anonymous):

y = x^2 - 10x + 21 this intercepts y axis at x = 0 substitute x = 0 in to get y coordinate to find the x intercept we need to solve for x x^2 - 10x + 21 = (x-7)(x-3) so when x^2 - 10x + 21 = 0 (x-7)(x-3) = 0 which happens at x = 7 or x = 3

OpenStudy (anonymous):

is y 21?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so y intercept is (0,21) x intercepts are (3,0) and (7,0)